Given: RS=ST                    .---------.---------.
Prove: RT=2ST                   R          S         T

What are the Statements and Reasons for this proof?

Answer by Edwin McCravy(20086)

Given: RS=ST .---------.---------.
Prove: RT=2ST R S T
What are the Statements and Reasons for this proof?

 Statements               Reasons

1. RT = RS + ST          1. A whole is equa1 to the sum of its parts.

2. RS = ST               2. Given

3. RT = ST + ST          3. A quantity may be substituted for its equal.
                           (Substitution postulate) (We are substituting
                           ST for RS in step 1.)

4. ST + ST = 2ST         4.    1+1=2

5. RT = 2ST              5  A quantity may be substituted for its equal.
                           (Substitution postulate) (We are substituting
                           2ST for ST + ST in step 3.)
-----------------------------------------------------------------

Edwin